# Overlap Gap and Computational Thresholds in the Square Wave Perceptron

### v2

@@ Abstract
Square Wave Perceptrons (SWPs) form a class of neural network models with oscillating activation function that exhibit intriguing “hardness” properties in the high-dimensional limit at a fixed signal-to-noise ratio $\alpha=O(1)$. In this work, we examine two key aspects of these models.
@@ 1 Introduction
Perceptron-based learning prototypes have been extensively studied by the statistical physics community, showing that despite their simplicity, which enables analytical investigations, these models exhibit complex behaviors, such as distinct thermodynamic phases separated by phase transitions. Two fundamental learning prototypes are the so-called “storage” and “teacher-student” problems. In both cases, the objective is to find a vector of weights $\boldsymbol{w}\in\{-1,+1\}^{N}$ s.t. the perceptron fits a database of $P$ given associations $\{(\boldsymbol{x}^{\mu},y^{\mu})\}_{\mu=1}^{P}\equiv\mathcal{D}$. In the statistical physics approach, this database is the source of “quenched disorder”. In other words, one requires that the following constraints are satisfied by $\boldsymbol{w}$ (see also Fig. 1):
@@
  • in the “teacher-student” (or planted) setting one draws a “teacher” set of weights, $\boldsymbol{w}^{T}\in\{-1,1\}^{N}$, with i.i.d. Rademacher elements $w^{T}_{i}$, $i\in[N]$, and the labels are given by:
    $y^{\mu}=\varphi\left(\frac{1}{\sqrt{N}}\boldsymbol{w}^{T}\cdot\boldsymbol{x}^{\mu}\right)\quad\forall\,\mu\in[P]\,.$ (2)
@@
Interestingly, the previous perceptron problems are not only studied as toy models for learning, but are also considered insightful examples of constraint satisfaction problems (CSPs) by various communities interested in aspects of algorithmic hardness. Examples range from the statistical physics of glasses [8] to theoretical computer science [9, 10]. In fact, depending on the fraction of the number of patterns $\alpha=P/N$, which plays the role of a control parameter, the behavior of polynomial-time algorithms designed to find a solution to constraints (1) could change dramatically. In particular, in the limit $P,N\rightarrow\infty$ held at constant ratio $\alpha$, there are regions of $\alpha$ where the constraints (1) are satisfiable w.h.p.  but no polynomial-time algorithm is known to find solutions. For the teacher-student problem, in cryptography parlance, the network is a one-way function (OWF) for these parameters. A fundamental open question concerns the mathematical characterization of these “hard” regions. From statistical physics investigations, the idea emerged that the hardness in CSPs can be understood in terms of changes of the geometrical properties of their solution space [11, 12, 13, 14, 15]. A disconnectivity transition, $\alpha_{U}$, predicted by Baldassi et al. (2015) [16], which they refer to as the local entropy transition, is found in the solution space of a binary perceptron problem. The transition $\alpha_{U}$ is interpreted as a separation between a regime ($\alpha<\alpha_{U}$) where the densest cores of solutions are immersed in a huge connected structure, and a regime ($\alpha_{U}<\alpha$) where this structure breaks up. Numerical studies of the performance of the best known polynomial solvers for the problem show that these algorithms stop working when $\alpha>\alpha_{U}$. Building upon this geometric point of view, [17] introduced the overlap gap property (OGP), which is a disconnectivity feature of the space of solution of CSPs related to the local entropy transition. Remarkably, the presence of OGP has been rigorously proven to imply the failure of a rather huge class of efficient algorithms, and it is conjectured to imply hardness against all efficient algorithms [18]  [^1] Constraint satisfaction problems with a linear structure like XOR-SAT are considered somehow an exception to the hardness conjecture, since despite the disconnectivity properties of their space of solutions, they can be always solved in polynomial time by Gaussian elimination..
In this paper, we study a family of oscillating activation functions of the form,
@@
that we call square wave perceptrons (SWPs). We show that SWPs exhibit extreme hardness properties in the limit of small $\delta$. We identify, both in the storage and in the teacher-student settings, the emergence of an overlap gap at a threshold, which can be made arbitrarily small by suitably increasing the frequency of oscillations $1/\delta$ of the activation function in Eq. 3. This suggests that in this small-$\delta$ regime, typical instances of the problem are hard to solve even for small values of $\alpha$. In the teacher-student setup, we also show that the recovery threshold of the planted signal for message-passing algorithms can be made arbitrarily large by reducing $\delta$, similarly to certain adversarially planted random boolean satisfiability problems [19].
These properties make SWPs both a challenging benchmark for algorithms and an interesting candidate for cryptographic applications [20].
The rest of the paper is organized as follows. In Section 2, we discuss the thermodynamic properties of perceptron problems, and we introduce the notion of “overlap gap” property. We also discuss the analytic framework for the study of the OGP in perceptron problems with generic activation functions. In Section 3, we study the storage problem for the SWP. We start by computing the satisfiability threshold and then move to the analytical determination of the $m$-OGP thresholds within an annealed and an RS ansatz for different values of $m$ and $\delta$. We also study the limit $\delta\rightarrow 0$: in this regime, we find that solutions exist up to $\alpha=1$, and all the region $10<\alpha<$ 1 presents an overlap gapped phase. In Section 4 we study the teacher-student problem for the SWP. We compute the threshold $\alpha_{T}(\delta)$ above which the planted solution becomes the unique solution to the problem. Similarly to the storage case, we compute the $m$-OGP thresholds within an annealed and an RS ansatz for different values of $m$ and $\delta$ and analyze the limit $\delta\rightarrow 0$. Finally, we analyze the algorithmic problem of recovering the planted signal above $\alpha_{T}(\delta)$.
@@ 2 Phase Transitions and the Overlap Gap Property
We call $\alpha_{\mathrm{OGP}}(m)$ the threshold above which $m-$OGP begins to hold. The $m-$OGP has a simple geometric interpretation. It requires that for every $m$-tuple of solutions, there are at least two solutions that are “close” or “far apart”. A technical remark is that to rule out stable algorithms, a contradiction argument is usually presented based on a more complicated “ensemble” variant of the $m$-OGP, which allows the configurations of the $m$-tuple in the previous definition to be solutions of possibly different instances of the problem. However, as argued in [6], the ensemble $m$-OGP and the non-ensemble $m$-OGP are found to take place at the exact same threshold. For this reason, we restrict our analysis to the non-ensemble version.
Note that if $m^{\prime}>m$, $m$-OGP implies $m^{\prime}$-OGP, and therefore $\alpha_{\mathrm{OGP}}(m^{\prime})\leq\alpha_{\mathrm{OGP}}(m)$. In particular, obstruction to stable algorithms does not require a specific value of $m$, implying that $\alpha_{\mathrm{OGP}}(m)$ represents an upper bound to the values of $\alpha$ that are accessible by stable algorithms, and the best upper bound is the one obtained in the limit $m\rightarrow\infty$ (that is taken after the limit $N\to\infty$).
@@ 2.1 The Clonated Free Entropy
$\mathbb{X}_{\mathcal{D}}(\boldsymbol{w})=\prod_{\mu=1}^{P}\Theta\left[\varphi\left(\frac{1}{\sqrt{N}}\boldsymbol{w}^{T}\cdot\boldsymbol{x}^{\mu}\right)\,\varphi\left(\frac{1}{\sqrt{N}}\boldsymbol{w}\cdot\boldsymbol{x}^{\mu}\right)\right]\,,$ (7)
enforces that all the clones are solutions of a teacher-student problem, with teacher $\boldsymbol{w}^{T}$. As we shall see, the storage setting can be obtained as a special case of the previous formulas. To shorten the notation, in the following we do not write explicitly the disorder dependence in the partition function.
The partition function (6) is a random variable that depends on the realization of the disorder. We are interested in studying the typical properties of the clonated system, i.e. the most probable value of (6), which in general does not coincide with its average. In the large $N$ limit, the most probable value $\hat{\mathcal{N}}_{m}(q)$ of $\mathcal{N}_{m}(q)$ is expected to be given by $\hat{\mathcal{N}}_{m}(q)\sim\exp{(m\,N\hat{\phi}_{m}(q))}$ [39], where $\hat{\phi}_{m}$ is the so-called quenched free entropy  [^2] For a proof of this statement in the ABP see [40].:
$\hat{\phi}_{m}(q)=\lim_{N\rightarrow\infty}\frac{1}{mN}\mathds{E}_{\mathcal{D}}\log{\mathcal{N}_{m}(q)}\,.$ (8)
@@
$\hat{\phi}_{m}(q)\leq\phi^{\mathrm{ann}}_{m}(q)=\lim_{N\to\infty}\frac{1}{mN}\ln\mathbb{E}_{\mathcal{D}}\mathcal{N}_{m}(q)\,.$ (9)
The upper bound follows from the fact that $\mathcal{N}_{m}(q)$ is a non-negative and integer valued random variable, therefore Markov inequality ensures that
$\text{Pr}_{\mathcal{D}}[\mathcal{N}_{m}(q)>0]\leq\mathbb{E}_{\mathcal{D}}\mathcal{N}_{m}(q)=\mathrm{e}^{mN\phi_{m}^{\mathrm{ann}}(q)}\,.$ (10)
@@
$p_{a}=\frac{1}{N}\sum_{i=1}^{N}w_{i}^{a}w_{i}^{T}\,.$ (12)
Assuming a symmetric ansatz for the overlaps, $p_{a}=p$ (see Section 4), and taking the large $N$ limit, $\phi_{m}^{\mathrm{ann}}(q)$ becomes:
@@
$\begin{split}G_{S}^{(m)}(q,p)&=\lim_{N\rightarrow\infty}\frac{1}{mN}\log{\sum_{\boldsymbol{w}^{1},\dots,\boldsymbol{w}^{m}}\prod_{a}\delta\left(p-\frac{\boldsymbol{w}^{T}\cdot\boldsymbol{w}}{N}\right)\,\prod_{a<b}\delta\left(q-\frac{\boldsymbol{w}^{a}\cdot\boldsymbol{w}^{b}}{N}\right)}=\\ &=\min_{\hat{q},\hat{p}}\left\{-\frac{(m-1)}{2}q\,\hat{q}-p\,\hat{p}-\frac{\hat{q}}{2}+\frac{1}{m}\log{\left(\sum_{k=0}^{m}\binom{m}{k}\mathrm{e}^{(2k-m)\hat{p}+\frac{\hat{q}}{2}(2k-m)^{2}}\right)}\right\}\,.\end{split}$ (15)
@@ 3 The Storage Problem
In this section, we focus on the storage problem. We start with the computation of the existence of solutions (critical capacity or SAT/UNSAT transition [43, 44, 29, 33, 45]) as a function of $\delta$, see Section 3.1. Then in Section 3.2, we discuss the presence of an OGP for $\alpha\geq\alpha_{\mathrm{OGP}}(\delta)$ as introduced in Section 2. In particular, we will see that in the small $\delta$ limit, solutions exist all the way up to $\alpha_{c}(0)=1$, but OGP holds for any $\alpha>0$.
@@ 3.1 Satisfiability Threshold
giving an upper bound estimate for the capacity $\alpha_{c}^{ann}(\delta)=1$ (see B.1). As shown in B, the replica symmetric (RS) ansatz for the free entropy of one clone $\phi^{rs}_{1}$ can be recovered from the annealed computation with $m$ clones (see Eq. (14)) as follows:
@@
In Fig. 2 we show the critical capacity estimated with the zero-entropy criterion on the RS free entropy, as a function of $\delta$. We note that this method [29] has been shown to be exact for ABP [46, 47], which corresponds to taking the SWP in the limit $\delta\rightarrow\infty$ [29]. We note that, in the small $\delta$ limit, the RS estimate of the capacity saturates the annealed upper bound. In C, using a second-moment argument, we show that the first moment estimate of the capacity is correct in this limit.
@@ 3.2 Overlap Gap
In Fig. 3(a), we show the annealed estimates of the $\alpha_{\mathrm{\mathrm{OGP}}}(m,\delta)$ thresholds for different values of $m$ and $\delta$. The curves are obtained by the numerical solution of Eqs. (11).
@@ 3.2.1 The Small $\delta$ Limit
The function $I(\bullet,\bullet)$ for the SWP reads (see Eq. (97)):
@@
By perturbatively solving Eqs. (11) in the small $\delta$ limit (see E) one finds that $\frac{1-q}{\delta^{2}}\to\infty$, implying that the series in (22) tends to zero term by term. Therefore, $G_{E}^{(m)}\to-\log 2$ and the free entropy has the following form:
@@
Since $G_{S}^{(m)}(q)$ is a monotonic decreasing function of $q$ (see D), the condition
@@
The expansion of Eqs. (11) presented in E provides a perturbative estimate $\widetilde{\alpha}^{ann}_{\mathrm{OGP}}(m,\delta)$ of $\alpha^{ann}_{\mathrm{OGP}}(m,\delta)$, which becomes accurate for small $\delta$. In E we note that $\widetilde{\alpha}^{ann}_{\mathrm{OGP}}(m,\delta)\geq\alpha^{ann}_{\mathrm{OGP}}(m,\delta)$, and therefore we can use $\min_{m}\widetilde{\alpha}^{ann}_{\mathrm{OGP}}(m,\delta)$ as an upper bound to the OGP threshold. In Fig. 3(a) we represent the limiting curve $\min_{m}\widetilde{\alpha}^{ann}_{\mathrm{OGP}}(m,\delta)$ as a function of $\delta$.

### v4

@@ Abstract
Square Wave Perceptrons (SWPs) form a class of neural network models with oscillating activation function that exhibit intriguing “hardness” properties in the high-dimensional limit at a fixed constraint density $\alpha=O(1)$. In this work, we examine two key aspects of these models.
@@ 1 Introduction
Perceptron-based learning prototypes have been extensively studied by the statistical physics community, showing that despite their simplicity, which enables analytical investigations, these models exhibit complex behaviors, such as distinct thermodynamic phases separated by phase transitions. Two fundamental learning prototypes are the so-called “storage” and “teacher-student” problems. In both cases, the objective is to find a vector of weights $\boldsymbol{w}\in\{-1,+1\}^{N}$ s.t. the perceptron fits a database of $P$ given associations $\{(\boldsymbol{x}^{\mu},y^{\mu})\}_{\mu=1}^{P}\equiv\mathcal{D}$. In the statistical physics approach, this database is the source of “quenched disorder”. In other words, one requires that the following constraints are satisfied by $\boldsymbol{w}$ (see also Fig.˜1):
@@
  • in the “teacher-student” (or planted) setting one draws a “teacher” set of weights, $\boldsymbol{w}^{*}\in\{-1,1\}^{N}$, with i.i.d. Rademacher elements $w^{*}_{i}$, $i\in[N]$, and the labels are given by:
    $y^{\mu}=\varphi\left(\frac{1}{\sqrt{N}}\boldsymbol{w}^{*}\cdot\boldsymbol{x}^{\mu}\right)\quad\forall\,\mu\in[P]\,.$ (2)
@@
Interestingly, the previous perceptron problems are not only studied as toy models for learning, but are also considered insightful examples of constraint satisfaction problems (CSPs) by various communities interested in aspects of algorithmic hardness. Examples range from the statistical physics of glasses [8] to theoretical computer science [9, 10]. In fact, depending on the fraction of the number of patterns $\alpha=P/N$, which plays the role of a control parameter, the behavior of polynomial-time algorithms designed to find a solution to constraints (1) could change dramatically. In particular, in the limit $P,N\rightarrow\infty$ held at constant ratio $\alpha$, there are regions of $\alpha$ where the constraints (1) are satisfiable w.h.p.  but no polynomial-time algorithm is known to find solutions. For the teacher-student problem, in cryptography parlance, the network is a one-way function (OWF) for these parameters. A fundamental open question concerns the mathematical characterization of these “hard” regions. From statistical physics investigations, the idea emerged that hardness in CSPs can be understood in terms of changes in the geometrical properties of their solution space [11, 12, 13, 14, 15]. A transition between a regime where the densest cores of solutions are immersed in a huge connected structure, and a regime where this structure breaks up is conjectured to mark the onset of algorithmic hardness. This transition, referred to as the local entropy transition, was found in the solution space of a binary perceptron problem [16], matching the threshold where the best known polynomial solvers for the problem stop working. Building upon this geometric point of view, [17] defined a disconnectivity feature of the space of solution of CSPs termed overlap gap property (OGP). Beyond capturing the topological features of the solution space emerging at the local entropy transition, the OGP — or stronger variants such as the ensemble-OGP — provably imply the failure of broad classes of efficient algorithms [18]. The observation that the best available efficient solvers for certain hard CSPs are hindered by the presence of an OGP has led to the conjecture that the OGP serves as a general predictor of algorithmic hardness across a wide range of problems, whose precise characterization remains an open question [^1] There exist problems where OGP alone is not predictive of hardness (see [19, 20, 21])..
In this paper, we study the onset of OGP and the consequent conjectured hardness thresholds for the teacher-student and storage problems in a family of oscillating activation functions of the form
@@
that we call square wave perceptrons (SWPs). We show that SWPs exhibit extreme hardness properties in the limit of small $\delta$. We identify, both in the storage and in the teacher-student settings, the emergence of an overlap gap at a threshold, which can be made arbitrarily small by suitably increasing the frequency of oscillations $1/\delta$ of the activation function in Eq.˜3. This suggests that in this small-$\delta$ regime, typical instances of the problem are hard to solve even for small values of $\alpha$. In the teacher-student setup, we also show that the recovery threshold of the planted signal for message-passing algorithms can be made arbitrarily large by reducing $\delta$, similarly to certain adversarially planted random boolean satisfiability problems [22]. In a recent line of research, the success of efficient algorithms has been linked to the existence of regions of densely packed solutions (see for example [23]). The intuition behind the hardness of the SWP as $\delta\to 0$, is that increasing the frequency of oscillations should destroy such dense entropic regions, as the labels become more and more sensitive to the weights.
These properties make SWPs both a challenging benchmark for algorithms and an interesting candidate for cryptographic applications. In particular, in [24] a design for Collision Resistant Hash Functions has been proposed, leveraging some aspects of computational hardness in the SWP. In short, Collision Resistant Hash Functions are functions for which finding two distinct inputs with the same output is computationally hard. As Collision Resistant Hash Functions are of fundamental importance, being at the basis of multiple cryptographic protocols and security guarantees [25, 26, 27, 28], it is interesting to explore other nuances of computational hardness stemming from SWPs activation in more canonical settings, as the storage and teacher-student problem.
The rest of the paper is organized as follows. In Section˜2, we discuss the thermodynamic properties of perceptron problems, and we introduce the notion of “overlap gap” property. We also discuss the analytic framework for the study of the OGP in perceptron problems with generic activation functions. In Section˜3, we study the storage problem for the SWP. We start by computing the satisfiability threshold and then move to the analytical determination of the $m$-OGP thresholds within an annealed and an RS ansatz for different values of $m$ and $\delta$. We also study the limit $\delta\rightarrow 0$: in this regime, we find that solutions exist up to $\alpha=1$, and all the region $0<\alpha<1$ presents an overlap gapped phase. In Section˜4 we study the teacher-student problem for the SWP. We compute the threshold $\alpha_{T}(\delta)$ above which the planted solution becomes the unique solution to the problem. Similarly to the storage case, we compute the $m$-OGP thresholds within an annealed and an RS ansatz for different values of $m$ and $\delta$ and analyze the limit $\delta\rightarrow 0$. Finally, we analyze the algorithmic problem of recovering the planted signal above $\alpha_{T}(\delta)$.
@@ 2 Phase Transitions and the Overlap Gap Property
We call $\alpha_{\mathrm{OGP}}(m)$ the threshold above which $m$-OGP begins to hold. The $m$-OGP has a simple geometric interpretation. It requires that for every $m$-tuple of solutions, there are at least two solutions that are “close” or “far apart”. A technical remark is that to rule out stable algorithms, a contradiction argument is usually presented based on a more complicated “ensemble” variant of the $m$-OGP, which allows the configurations of the $m$-tuple in the previous definition to be solutions of possibly different instances of the problem. However, as argued in [6], the ensemble $m$-OGP and the non-ensemble $m$-OGP are often found to take place at the exact same threshold (we mention here as an example the case of finding the largest independent set in sparse Erdos Renyi random graphs or the symmetric perceptron problem). We refer the reader to B for additional details on the non-ensemble version and why we expect that it coincides to OGP in our case. For this reason, from now on we will restrict our analysis to the non-ensemble version.
Note that if $^{\prime}>m, mm$-OGP implies $m^{\prime}$-OGP, and therefore $\alpha_{\mathrm{OGP}}(m^{\prime})\leq\alpha_{\mathrm{OGP}}(m)$. In particular, obstruction to stable algorithms does not require a specific value of $m$, implying that $\alpha_{\mathrm{OGP}}(m)$ represents an upper bound to the values of $\alpha$ that are accessible by stable algorithms, and the best upper bound is the one obtained in the limit $m\rightarrow\infty$ (that is taken after the limit $N\to\infty$).
@@ 2.1 The Clonated Free Entropy
$\mathbb{X}_{\mathcal{D}}(\boldsymbol{w})=\prod_{\mu=1}^{P}\Theta\left[\varphi\left(\frac{1}{\sqrt{N}}\boldsymbol{w}^{*}\cdot\boldsymbol{x}^{\mu}\right)\,\varphi\left(\frac{1}{\sqrt{N}}\boldsymbol{w}\cdot\boldsymbol{x}^{\mu}\right)\right]\,$, (7)
where we denote by $\Theta(\bullet)$ the Heaviside step function, enforces that all clones are solutions of a teacher-student problem, with teacher $\boldsymbol{w}^{*}$. As we shall see, the storage setting can be obtained as a special case of the previous formulas. To shorten the notation, in the following we do not write explicitly the disorder dependence in the partition function.
The partition function (6) is a random variable that depends on the realization of the disorder. We are interested in studying the typical properties of the clonated system, i.e. the most probable value of (6), which in general does not coincide with its average. In the large $N$ limit, the most probable value $\hat{\mathcal{N}}_{m}(q)$ of $\mathcal{N}_{m}(q;\mathcal{D})$ is expected to be given by $\hat{\mathcal{N}}_{m}(q)\sim\exp{(m\,N\hat{\phi}_{m}(q))}$ [45], where $\hat{\phi}_{m}$ is the so-called quenched free entropy  [^2] For a proof of this statement in the ABP see [46].:
$\hat{\phi}_{m}(q)=\lim_{N\rightarrow\infty}\frac{1}{mN}\mathds{E}_{\mathcal{D}}\log{\mathcal{N}_{m}(q;\mathcal{D})}\,.$ (8)
@@
$\hat{\phi}_{m}(q)\leq\phi^{\mathrm{ann}}_{m}(q)=\lim_{N\to\infty}\frac{1}{mN}\ln\mathbb{E}_{\mathcal{D}}\mathcal{N}_{m}(q;\mathcal{D})\,.$ (9)
The upper bound follows from the fact that $\mathcal{N}_{m}(q;\mathcal{D})$ is a non-negative and integer valued random variable, therefore Markov inequality ensures that
$\text{Pr}_{\mathcal{D}}[\mathcal{N}_{m}(q;\mathcal{D})>0]\leq\mathbb{E}_{\mathcal{D}}\mathcal{N}_{m}(q;\mathcal{D})=\mathrm{e}^{mN\phi_{m}^{\mathrm{ann}}(q)}\,.$ (10)
@@
$p_{a}=\frac{1}{N}\sum_{i=1}^{N}w_{i}^{a}w_{i}^{*}\,.$ (12)
Assuming a symmetric ansatz for the overlaps, $p_{a}=p$ (see Section˜4), and taking the large $N$ limit, $\phi_{m}^{\mathrm{ann}}(q)$ becomes:
@@
[…]
@@ 3 The Storage Problem
In this section, we focus on the storage problem. We start with the computation of the existence of solutions (critical capacity or SAT/UNSAT transition [49, 50, 36, 23, 51]) as a function of $\delta$, see Section˜3.1. Then in Section˜3.2, we discuss the presence of an OGP for $\alpha\geq\alpha_{\mathrm{OGP}}(\delta)$ as introduced in Section˜2. In particular, we will see that in the small $\delta$ limit, solutions exist all the way up to $\alpha_{c}(0)=1$, but OGP holds for any $\alpha>0$.
@@ 3.1 Satisfiability Threshold
giving an upper bound estimate for the capacity $\alpha_{c}^{ann}(\delta)=1$ (see C.1). As shown in C, the replica symmetric (RS) ansatz for the free entropy of one clone $\phi^{rs}_{1}$ can be recovered from the annealed computation with $m$ clones (see Eq. (14)) as follows:
@@
In Fig.˜2 we show the critical capacity estimated with the zero-entropy criterion on the RS free entropy, as a function of $\delta$. We note that this method [36] has been shown to be exact for ABP [52, 53], which corresponds to taking the SWP in the limit $\delta\rightarrow\infty$ [36]. We note that, in the small $\delta$ limit, the RS estimate of the capacity saturates the annealed upper bound. In D, using a second-moment argument, we show that the first moment estimate of the capacity is correct in this limit.
@@ 3.2 Overlap Gap
In Fig.˜3(a), we show the annealed estimates of the $\alpha_{\mathrm{\mathrm{OGP}}}(m,\delta)$ thresholds for different values of $m$ and $\delta$. The curves are obtained by the numerical solution of Eqs. (11).
@@ 3.2.1 The Small $\delta$ Limit
The function $I(\bullet,\bullet)$ depends on the specific activation function. In the case of the SWP, we denote it by $I_{\delta}(\bullet,\bullet)$ (see Eq. (99)), with the explicit form:
@@
By perturbatively solving Eqs. (11) in the small $\delta$ limit (see G) one finds that $\frac{1-q}{\delta^{2}}\to\infty$, implying that the series in (22) tends to zero term by term. Therefore, $G_{E}^{(m)}\to-\log 2$ and the free entropy has the following form:
@@
Since $G_{S}^{(m)}(q)$ is a monotonic decreasing function of $q$ (see F), the condition
@@
The expansion of Eqs. (11) presented in G provides a perturbative estimate $\widetilde{\alpha}^{ann}_{\mathrm{OGP}}(m,\delta)$ of $\alpha^{ann}_{\mathrm{OGP}}(m,\delta)$, which becomes accurate for small $\delta$. In G we note that $\widetilde{\alpha}^{ann}_{\mathrm{OGP}}(m,\delta)\geq\alpha^{ann}_{\mathrm{OGP}}(m,\delta)$, and therefore we can use $\min_{m}\widetilde{\alpha}^{ann}_{\mathrm{OGP}}(m,\delta)$ as an upper bound to the OGP threshold. In Fig.˜3(a) we represent the limiting curve $\min_{m}\widetilde{\alpha}^{ann}_{\mathrm{OGP}}(m,\delta)$ as a function of $\delta$.